OpenStudy (anonymous):

Write and equation for each line. Through (2,2) y-intercept 10

OpenStudy (anonymous):

general formula of a straight line is y=ax+b b is the y-intercept y=ax+10 substitute the point (2,2) into the equation to find a 2=a(2) + 10 2=2a + 10 2-10=2a -8=2a a=-8/2 a=-4 substitute the value of a=-4 back into the original equation with the y-intercept y=-4x+10, this is your final answer.

OpenStudy (anonymous):

What if you had to find y then what would you do?

OpenStudy (anonymous):

then they most likely would tell you the gradient, y-intercept and when x=3 find y

OpenStudy (anonymous):

or something similar, it really dependso n the question.

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

no problem, medal thanks :)